OpenStudy (unklerhaukus):

For the following differential equations, check that \(y_1\) is a solution of the corresponding homogeneous equation and then calculate the general solution by the reduction of order method. \[\left(1+x^2\right)y^{\prime\prime}+xy^\prime-y=0;\qquad y_1=x\]

OpenStudy (wasiqss):

eqaution is not visible

OpenStudy (unklerhaukus):

\[\left(1+x^2\right)y^{\prime\prime}+xy^\prime-y=0;\qquad y_1=x\]\[\qquad\qquad\qquad\qquad\qquad\qquad y_1^\prime =1\]\[\qquad\qquad\qquad\qquad\qquad\qquad y_1^{\prime\prime} =0\] \[\left(1+x^2\right)(0)+x(1)-(x)=0\] \[\therefore\quad y_1=x \qquad\text{is a solution}\] \[y_2=uy_1=ux\]\[y_2^\prime=u^\prime x+u\]\[y_2^{\prime\prime}=u^{\prime\prime}x+u^\prime+u^\prime=u^{\prime\prime}x+2u^\prime\]\[\left(1+x^2\right)y^{\prime\prime}_2+xy^\prime_2-y_2=0\]\[\left(1+x^2\right)\left(u^{\prime\prime}x+2u^\prime\right)+x\left(u^\prime x+u\right)-\left(ux\right)=0\]\[\left(1+x^2\right)u^{\prime\prime}x+\left(2+2x^2+x^2\right)u^\prime+\left(x-x\right)u=0\]\[\left(x+x^3\right)u^{\prime\prime}+\left(2+3x^2\right)u^\prime=0\]

OpenStudy (unklerhaukus):

am i doing this right , i think i have to solve for \(u\)

OpenStudy (unklerhaukus):

\[u(x)\]

OpenStudy (wasiqss):

you gonnna solve by frobenius?

OpenStudy (anonymous):

@UnkleRhaukus ur doing right solve for u

OpenStudy (anonymous):

let u'=p

OpenStudy (unklerhaukus):

ah \[\left(x+x^3\right)u^{\prime\prime}+\left(2+3x^2\right)u^\prime=0\] is a \(u\)- absent equation

OpenStudy (anonymous):

yes thats right

OpenStudy (unklerhaukus):

\[p=u^\prime\]\[p^\prime=u^{\prime\prime}\]

OpenStudy (unklerhaukus):

\[\left(x+x^3\right)p^{\prime}+\left(2+3x^2\right)p=0\]\[p^{\prime}+\frac{2+3x^2}{x+x^3}p=0\]

OpenStudy (unklerhaukus):

\[\large{\mu(x)=e^{\int{\frac{2+3x^2}{x+x^3}}\text dx}}\]

OpenStudy (unklerhaukus):

is this a easy integral ?,

OpenStudy (anonymous):

just notice \[\frac{2+3x^2}{x+x^3}=\frac{1}{x+x^3}+\frac{1+3x^2}{x+x^3}=\frac{1}{x}-\frac{1}{2}\frac{2x}{1+x^2}+\frac{1+3x^2}{x+x^3}\]

OpenStudy (unklerhaukus):

you have used partial fractions/

OpenStudy (anonymous):

yes it is

OpenStudy (unklerhaukus):

ill work on getting that decomposition by myself, ill be back when i have the details

OpenStudy (anonymous):

ok

OpenStudy (unklerhaukus):

\[\frac{2+3x^2}{x\left(x+x^2\right)}=\frac{1}{x\left(x+x^2\right)}+\frac{1+3x^2}{x\left(x+x^2\right)}\] \[\frac{1}{x\left(x+x^2\right)}=\frac Ax+\frac B{x(1+x^2)}\] \[1=x(1+x^2)A+Bx\]

OpenStudy (unklerhaukus):

im not how to can find \(A,B\)

OpenStudy (unklerhaukus):

* i am not sure how to find \(A,B\)

OpenStudy (anonymous):

\[we \ \ have \ \ \frac{2+3x^2}{x+x^3} \ \ \ \ \] not \[\frac{2+3x^2}{x(x+x^3)}\]

OpenStudy (unklerhaukus):

i reverted back to \[x+x^3=x(1+x^2)\] to make the partial fractions decomposition easier

OpenStudy (unklerhaukus):

i see i have made some mistake s

OpenStudy (unklerhaukus):

\[x\left(1+x^2\right)u^{\prime\prime}+\left(2+3x^2\right)u^\prime=0\]\[x\left(1+x^2\right)p^{\prime}+\left(2+3x^2\right)p=0\]\[p^{\prime}+\frac{2+3x^2}{x\left(1+x^2\right)}p=0\]\[\large{\mu(x)=e^{\int{\frac{2+3x^2}{x\left(1+x^2\right)}}\text dx}}\]

OpenStudy (unklerhaukus):

\[\frac{2+3x^2}{x\left(1+x^2\right)}=\frac{1}{x\left(1+x^2\right)}+\frac{1+3x^2}{x\left(x+x^2\right)}\]\[\frac{1}{x\left(1+x^2\right)}=\frac Ax+\frac B{x(1+x^2)}\]\[1=x(1+x^2)A+Bx\]

OpenStudy (anonymous):

use this \[\frac{A}{x}+\frac{Bx+C}{1+x^2}\]

OpenStudy (unklerhaukus):

ah clearly my memory is terrible

OpenStudy (unklerhaukus):

\[1=x(1+x^2)A+{(Bx+C)}x\]\[1=Ax+Ax^3+Bx^2+Cx\]\[1=Ax^3+Bx^2+(A+C)x\]

OpenStudy (unklerhaukus):

how do i find the constants?

OpenStudy (anonymous):

i think u made a little mistake in multiply \[1=(A+B)x^2+Cx+A\]

OpenStudy (unklerhaukus):

\[\frac{1}{x\left(1+x^2\right)}=\frac Ax+\frac {Bx+C}{1+x^2}\]\[1=(1+x^2)A+{(Bx+C)}x\]\[1=A+Ax^2+Bx^2+Cx\]\[1=A+(A+B)x^2+Cx\]

OpenStudy (unklerhaukus):

so A=1 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

B=-1 and C=0

OpenStudy (unklerhaukus):

how do you find those

OpenStudy (anonymous):

A+B=0 and C=0

OpenStudy (anonymous):

think of \[1=0*x^2+0*x+1\]

OpenStudy (unklerhaukus):

\[1=A+(A+B)x^2+Cx\] \[A=1\]\[(A+B)=0;\qquad B=0\]\[C=0\]

OpenStudy (unklerhaukus):

** B=-1

OpenStudy (unklerhaukus):

\[\frac{1}{x\left(1+x^2\right)}=\frac Ax+\frac {Bx+C}{1+x^2}=\frac 1x-\frac {x}{1+x^2}\]

OpenStudy (anonymous):

exactly

OpenStudy (unklerhaukus):

\[\frac{1+3x^2}{x\left(1+x^2\right)}=\frac{1}{x(1+x^2)}+\frac{3x}{1+x^2}\]\[\frac{1}{x(1+x^2)}=\frac Dx+\frac {Ex+F}{1+x^2} \]\[1=D(1+x^2)+(Ex+F)x\]\[1=D+(D+E)x^2+Fx\] \[D=1\]\[(D+E)=0;\qquad E=-1\]\[F=0\] \[\frac{1}{x(1+x^2)}=\frac 1x-\frac {x}{1+x^2} \]

OpenStudy (anonymous):

\[\int\limits \frac{2+3x^2}{x+x^3}dx=\int\limits (\frac{1}{x}-\frac{1}{2}\frac{2x}{1+x^2}+\frac{1+3x^2}{x+x^3})dx\]

OpenStudy (anonymous):

its not Necessary to fractionate \[\frac{1+3x^2}{x+x^3}\] because \[\frac{d}{dx}(x+x^3)=1+3x^2\]

OpenStudy (unklerhaukus):

damn your right. that saves a lot of work (That i have already done) oh well good practice .

OpenStudy (experimentx):

\[ (y'(1+x^2))' = (yx)' \\ y'(1+x^2) = yx + c\\ \]

OpenStudy (experimentx):

now that's a linear first order DE

OpenStudy (anonymous):

thats right but we want to use the reduction of order method

OpenStudy (unklerhaukus):

\[\frac{2+3x^2}{x\left(1+x^2\right)}=\frac 1x-\frac {x}{1+x^2}+\frac{1+3x^2}{x\left(1+x^2\right)}\] \[\frac{2+3x^2}{x\left(1+x^2\right)}=\frac 1x-\frac 12\frac {2x}{1+x^2}+\frac{1+3x^2}{x\left(1+x^2\right)}\] \[\large{\mu(x)=e^{\int\frac 1x-\frac 12\frac {2x}{1+x^2}+\frac{1+3x^2}{x\left(1+x^2\right)}\text dx}}\]

OpenStudy (experimentx):

isn't that the reduction of order??

OpenStudy (anonymous):

take a look http://en.wikipedia.org/wiki/Reduction_of_order

OpenStudy (unklerhaukus):

\[\large{\mu(x)=e^{\ln|x|-\ln|1+x^2|+\ln\left|x(1+x^2)\right|}}\]\[=e^{\ln\left|\frac{x^2(1+x^2)}{1+x^2}\right|}\]\[\mu(x)=x^2\]

OpenStudy (anonymous):

u forgotted 1/2 check the integral again

OpenStudy (unklerhaukus):

your right.

OpenStudy (unklerhaukus):

\[\large{\mu(x)=e^{\ln|x|-\ln\sqrt{1+x^2}+\ln\left|x(1+x^2)\right|}}\] \[=e^{\ln\left|\frac{x^2(1+x^2)}{\sqrt{1+x^2}}\right|}\] \[\mu(x)=x^2\sqrt{1+x^2}\]

OpenStudy (unklerhaukus):

\[\left(x^2\sqrt{1+x^2} p\right)^{\prime}=0\]\[x^2\sqrt{1+x^2}p=1\]\[p=\frac{1}{x^2\sqrt{1+x^2}}\]\[u^{\prime}=\frac{1}{x^2\sqrt{1+x^2}}\]

OpenStudy (unklerhaukus):

\[u=\int\frac{1}{x^2\sqrt{1+x^2}}\text dx\] do i have to use partial fractions decomposition again now///

OpenStudy (anonymous):

no

OpenStudy (unklerhaukus):

have i made another bunch of mistakes? or is integration by parts a better idea? there are integral similar to this in my standard table of [41] integrals but i cant see any that match

OpenStudy (unklerhaukus):

wolfram is suggesting a trig substitution,

OpenStudy (anonymous):

sorry i lost my connection about hour ago

OpenStudy (anonymous):

yes x=tan t

OpenStudy (unklerhaukus):

i dont really like trig substitution because i dont understand why they are valid,

OpenStudy (anonymous):

me too

OpenStudy (unklerhaukus):

\[u=-\frac{\sqrt{x^2+1}}{x}\]\[y_2=uy_1=-{\sqrt{x^2+1}}\] \[y=ay_1+by_2\]\[y= ax+b{\sqrt{x^2+1}}\]